A) How many milliliters of 0.145 M HCI are needed to neutralize completely 35.0 mL of 0.101 M Ba(OH)2 solution?

B) How many milliliters of 2.50 M H2SO4 are needed to neutralize 50.0g of NaOH?

C) If 56.8 mL of BaCl2 solution is needed to precipitate all the sulfate in a 534 mg sample of Na2SO4 (forming BaSO4), what is the molarity of the solution?

D) If 37.5 mL of 0.375 M HCI solution is needed to neutralize a solution of Ca(OH)2, how many grams of Ca(OH)2 must be in the solution?


1
Expert's answer
2021-02-26T06:01:18-0500

A) 0.035*0.101/0.145 =0,02438 L.

B) 50/40*2*2.5=0.25L = 250 ml.

C) 0.534/142*0.0568=0,0662 M

D) 0.0375*0.375*74=1,040625 g.